19.5g CH3COOH is dissolved in 500g of water the depression in the freezing point observed is 1.0k calculate van't hoff factor and dissociation constant of the acid.
1
Expert's answer
2020-04-28T12:25:04-0400

‎ΔTfreezing=iCmK=1K

Cm=m( CH3COOH)/(Mr( CH3COOH)*m(H2O))=19.5/(60*0.5)=0.65 mol/kg

i=1/(1.86*0.65)=0.83;

Kd=(1*0.5/1.86)/(19.5/60)=0.83

Answer:i=0.83;

Kd=0.83